OpenStudy (anonymous):

Hey plz help (X-4)^3 I know the cube rule (X-4) (x^2+4x+16) I have to expand and simplify so I am getting X^3-64 But the answer is X^3-12x^2+48x-64 I think i dunno some rule

OpenStudy (anonymous):

I think that when multiplying your doing a sign error. Like for example your replacing a minus with a plus and things like that. A couple of reminders: -(x+6) = -x - 6, (-x)(-y) = xy (x)(y) = xy (-x)(y) = (-xy) (x)(-y) = (-xy) I used to do these mistakes so I totally understand what your currently going though. Hope this helps.

OpenStudy (anonymous):

@Boblovesmath thank u but i tried again it didnt help (X-4)(x^2+4x+16) X^3+4x^2+16x-4x^2-16x-64 X^3-64

OpenStudy (anonymous):

@dan815 can u plz help me i am sure u know it only for a sec plz

OpenStudy (phi):

do you know the distributive property? a(b+c)= ab + ac you can use that idea with your problem (x-4) (x^2+4x+16) distribute the (x-4) (multiply each term inside the parens by (x-4) ) what do you get ?

OpenStudy (anonymous):

X^3 -64

OpenStudy (anonymous):

See above^

OpenStudy (phi):

no. Just write down (x-4) times each term inside the parens do not simplify (yet) what do you get ?

OpenStudy (anonymous):

U mean (x-4)x^2 +(x-4)4x+(x-4)16

OpenStudy (phi):

yes, looks good.

OpenStudy (phi):

now notice for the first term that you wrote (x-4)x^2 or x^2(x-4) we can distribute the x^2 can you do that ?

OpenStudy (anonymous):

X^3-4x^2

OpenStudy (phi):

ok, now do the 2nd term (x-4)4x (we can distribute the 4x)

OpenStudy (anonymous):

4x^2-16

OpenStudy (phi):

and the last term (x-4)16

OpenStudy (anonymous):

16x-64

OpenStudy (phi):

X^3-4x^2 + 4x^2-16 +16x -64

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

4x^2-16x*

OpenStudy (phi):

I think we would get the correct answer if we started with (x-4)(x-4) = x(x-4) + -4(x-4) = x^2 -4x -4x +16 = x^2 -8x +16 now do (x-4)( x^2 -8x +16)

OpenStudy (anonymous):

(X-4) ^3 is cube why u put only two brackets of (x-4)

OpenStudy (phi):

It's easier to see what I did if we say A(x-4) = xA + -4A and then say, A is short for the first (x-4) x(x-4) + -4(x-4)

OpenStudy (phi):

**only two brackets of (x-4) because I am doing it in steps: (x-4) * ( (x-4)(x-4) ) the (x-4)(x-4) = x^2 -8x +16 I redid that because up top you wrote *** I know the cube rule (X-4) (x^2+4x+16) *** and that is not correct.

OpenStudy (phi):

if we use the correct (x-4)(x-4) = x^2 -8x +16 we can now do (x-4)(x^2 -8x +16) and it will turn out to be \[ x^3-4x^2 -8x^2 +32x +16x -64 \\ x^3 -12x^2 +48x -64 \]

OpenStudy (anonymous):

Oh thank u thank u